Capacitors of 2μF and 6μF are connected to a 100V battery. How much charge I supplied by the battery in charging the capacitors when (a) they are connected in series and (b) when they are connected in parallel?


1
Expert's answer
2022-03-02T10:49:10-0500

(a)

"C_{eq}=\frac{C_1C_2}{C_1+C_2}=\frac{2*6}{2+6}=1.5\:\rm \mu F"

"q_1=q_2=C_{eq}V\\
=1.5*10^{-6}*100=1.4*10^{-4}\:\rm C"

(b)

"q_1=C_{1}V\\
=2*10^{-6}*100=2*10^{-4}\:\rm C"

"q_2=C_{2}V\\
=6*10^{-6}*100=6*10^{-4}\:\rm C"